OpenStudy (anonymous):

\[Cosec ^{2}(\frac{ 5 }{ 6 }\pi)\]

OpenStudy (anonymous):

\[cosec ^{2}(\pi - \frac{ \pi }{ 6 })\] \[cosec ^{2} \frac{ \pi }{ 6 }\] =4